OpenStudy (anonymous):

Medal! Math help! -3 x + 7 =37

OpenStudy (yese30):

37-7=30 30 divide -3 = -10

OpenStudy (yese30):

does this help?

OpenStudy (anonymous):

no

OpenStudy (michele_laino):

If I subtract \(7\) from both sides, I get: \(-3x+7-7=37-7\) please simplify

OpenStudy (anonymous):

0 and 30

OpenStudy (michele_laino):

ok! I get: \(-3x=30\) Now i divide both sides by \(-3\) so i get: \[\frac{{ - 3x}}{{ - 3}} = \frac{{30}}{{ - 3}}\] again, please simplify

OpenStudy (anonymous):

0 and 30 i would show you how i answered but my phone is dead so i cant take a picture so i can post it

OpenStudy (anonymous):

help me

OpenStudy (michele_laino):

hint: \((-3)/(-3)=1\)

OpenStudy (anonymous):

1 and 10

OpenStudy (anonymous):

@dawsonflores how rude

OpenStudy (michele_laino):

we have: \(30/ (-3)=-10\)

OpenStudy (anonymous):

soo its -10

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

thank you baby doll